Twice a number minus a second number is -1. Twice the second number added. To three time the first number is 9. What are the two numbers
Answer by Seutip(231)

The equations are:
Let x- be the number
y- be another number
(1) 2x-y=-1
(2) 2y+3x=9
(1) 2x-y=-1
-y=-1-2x
y=1+2x
(2) 2y+3x=9
2(1+2x)+3x=9
2+4x+3x=9
7x=7 Therefore,
x=1
y=1+2x
y=1+2(1)
y=3
HOPETHATHELPS! <3
Follow me @seutip in instagram and twitter! <3